### v1

# Logic Augmented Generation

Abstract
Semantic Knowledge Graphs (SKG) face challenges with scalability, flexibility, contextual understanding, and handling unstructured or ambiguous information. However, they offer formal and structured knowledge enabling highly interpretable and reliable results by means of reasoning and querying. Large Language Models (LLMs) overcome those limitations making them suitable in open-ended tasks and unstructured environments. Nevertheless, LLMs are neither interpretable nor reliable. To solve the dichotomy between LLMs and SKGs we envision Logic Augmented Generation (LAG) that combines the benefits of the two worlds. LAG uses LLMs as Reactive Continuous Knowledge Graphs that can generate potentially infinite relations and tacit knowledge on-demand. SKGs are key for injecting a discrete heuristic dimension with clear logical and factual boundaries. We exemplify LAG in two tasks of collective intelligence, i.e., medical diagnostics and climate projections. Understanding the properties and limitations of LAG, which are still mostly unknown, is of utmost importance for enabling a variety of tasks involving tacit knowledge in order to provide interpretable and effective results.
keywords:
Knowledge Graphs , Large Language Models , Logic Augmented Generation
journal: Journal of Web Semantics
[…]
1 Introduction
Semantic Knowledge Graphs (SKGs) provide structured, interpretable knowledge and logical querying capabilities, making them valuable tools for modelling rich, linked datasets in diverse domains. However, their limitations in scalability, flexibility, and the integration of non-standard input restrict their application in open-ended tasks requiring collective intelligence, such as medical diagnostics and climate services. To address these challenges, we introduce Logic Augmented Generation (LAG), a novel paradigm that integrates SKGs with Large Language Models (LLMs) to leverage the strengths of both approaches. LAG conceptualises LLMs as potential Reactive Continuous Knowledge Graphs (RCKGs), which can dynamically adapt to diverse inputs by extending and contextualising SKGs that work as base models. SKGs ensure logical consistency, enforce factual boundaries, and foster interoperability, while LLMs process unstructured data and provide contextual insights and tacit knowledge on demand. This dual-layered approach enhances interpretability and reliability, mitigating the lack of truth-theoretic semantics in LLM’s output. We expect that LAG methods enable effective collaboration among experts and support the co-creation of actionable insights in complex, evolving contexts such as climate adaptation and collective medical decision-making.
2 Use case and problem framing
Semantic Knowledge Graphs (SKGs) provide structured, interpretable knowledge and support precise, logical querying. Accordingly they have been used for modelling rich linked datasets in a variety of domains, such as health, cultural heritage, social and climate sciences, etc. In the context of the HACID [^3] The HACID acronym stands for Hybrid Human Artificial Collective Intelligence in Open-Ended Decision Making project. project [1] we designed and modelled a SKG to support experts in open-ended diagnostic tasks with a collective intelligence approach in two different domains, i.e., medical diagnostics and climate services. Collective intelligence is the shared or group intelligence that emerges from the collaboration, collective efforts, and interactions of a group of individuals or entities. This concept is rooted in the idea that a group, whose members work together effectively, can achieve insights, solve problems, or make decisions that are beyond the capability of a single individual. Our SKG integrates different existing knowledge bases and provides a coherent conceptualisation formalised in a rich modular ontology network designed by reusing Ontology Design Patterns [2] (ODPs) and leveraging DOLCE-Zero [3][4] as a top-level ontology. For medical diagnostics we integrated data gathered from SNOMED-CT, ICD-10, and Wikidata.For climate services we integrated data from the Coupled Model Intercomparison Project [^4] https://www.wcrp-climate.org/wgcm-cmip. (CMIP) [5]. For both use cases the objective is to leverage the SKG for smoothening the collaboration and communication among experts by harmonising the diversity of thought in a context in which no single individual has all the answers, but the group as a whole possesses the knowledge needed to solve a problem. In medical diagnostics, the ultimate goal is to identify the correct diagnosis for a patient, thereby reducing errors, preserving life, and minimising the costs associated with incorrect treatments. Climate services urgently require systems that leverage collective intelligence to navigate and integrate vast, diverse datasets. By fostering collaboration among scientists, stakeholders, and AI-driven tools, challenges in the interpretability of results and the handling of unstructured data can be addressed more effectively. Nevertheless, the effectiveness of a SKG solely based on symbolic knowledge representation and reasoning remains limited when we need to support open-ended tasks with expansive problem spaces, such as collective intelligence-based medical diagnostics, or climate services. This limitation arises from the challenges of harmonizing and integrating standardised knowledge in the reference SKG with non-standard, natural language inputs from multiple experts. Firstly, the different inputs provided by experts might express different viewpoints in different languages and factual granularities. Secondly, tacit and domain-specific background knowledge might be required in order to make sense of those inputs, and making a superior collective knowledge emerge. Moreover, SKGs struggle with scalability, flexibility, evolvability, and the integration of complex, context-rich information. E.g., climate data include real-time environmental observations, policy impacts, and social adaptation strategies. Large Language Models (LLMs), in contrast, are adept at managing unstructured data and generating novel insights on demand, making them well-suited for open-ended tasks. Yet, their lack of interpretability and reliability hinders their application in fields where precise and accountable outputs are critical.
3 LLMs as Reactive Continuous Knowledge Graphs with logic boundaries
To address this gap, we use LLMs as potential Reactive Continuous Knowledge Graphs (RCKGs). RCKGs are SKGs that are extracted from possibly multimodal signal (e.g. texts, speech, pictures, sensory data, etc.) using (typically or mostly) continuous vector spaces such as generative pre-trained transformers. RCKG extraction happens by either mapping signal to natural language, which is supramodal  [^5] Natural language use by humans is based on synthesis and integration across senses: it depends on the multiple perception and action modalities of the users: it is supramodal. Knowledge Graphs can be considered amodal because they do not depend on any specific modality: they abstract beyond sensory input., and then extracting an amodal SKG from natural language, or directly extracting a KG from signal.
A SKG can be generated from text in multiple ways. State-of-the-art SKG extraction from text assumes situations as occurrences of conceptual frames, following the knowledge extraction paradigm introduced by the FRED machine reader [6], which extracts OWL SKGs based on Framester Semantics [7], and is also used [8] to formalise Abstract Meaning Representation [9] (AMR) graphs. Once extracted, a SKG is passed (within an augmented prompt) to an LLM as a source of extended meaning, to contextualise and complete a SKG with implicit and local knowledge. In this sense, a LLM is used as a reactive AI, which generates a piece of extended knowledge $K_{ext}$ when stimulated by an input piece of knowledge $K_{inp}$.
A first implementation of LAG as SKG-augmented RCKG generation is demoed as a multimodal knowledge extractor [^6] https://arco.istc.cnr.it/itaf/ [10], used to generate the example graph in Figure 2.
In addition to this definition of reactive and (mathematically) continuous KGs, Continuous Knowledge Graphs have been introduced recently, e.g. in [11] as evolutionary dynamic graphs, featuring incremental knowledge updates, schema adaptation, conflict resolution, source tracking and verification, and temporal awareness. The authors propose to reduce extraction to the normalization and manipulation of elementary propositions, but recognize that this is a very challenging task. As said above, logic augmented generation is better suited to represent situations as occurrences of conceptual frames in a frame semantics-compatible logic like OWL2.
The notion of continuity proposed by [11] focuses on the process of updating dynamic and adaptable knowledge representation systems that can evolve and integrate new information systematically over time. They talk of the well-known task of making graphs evolve continuously over time and available information. RCKGs are continuous firstly because LLMs are reactive continuous models: they are mostly trained on a finite and discrete corpus of data, which forms the basis of their learned knowledge; however, due to their ability to generalise from training data, they can generate a potentially infinite number of outputs by combining concepts, patterns and relationships in novel ways. This generative capability is what allows LLMs to learn and generate responses based on context-sensitive input data (such as text) and to continuously adapt to new contexts or queries during in-context learning. Hence, by means of prompting, their internal representation of knowledge adapts dynamically as more data is processed, allowing them to dynamically adjust and generate relevant responses based on the latest input. Accordingly, if properly prompted to genrate a KG, a LLM can generate potentially infinite triples by means of in-context learning. This means, for example, that LLMs do not require retraining, adjustments or evolution to their underlying data to generate new knowledge. Instead, by leveraging their ability to adapt to specific prompts, LLMs can dynamically create knowledge representation based on current inputs, making them a flexible tool for continuous knowledge generation without modifying the original knowledge corpus.
4 LAG: SKG+RCKG
RCKGs are an example of a neuro-symbolic approach that generates a much richer and deeper knowledge graph than the basic meaning of the original text. RCKGs extract tacit knowledge by capturing insights, reasoning, or behaviours that resemble the informal, experiential, and often unspoken knowledge humans acquire through experience and practice [12]. While tacit knowledge is difficult to articulate or codify, it plays a crucial role in problem-solving and decision-making, making it particularly valuable for collective intelligence tasks that require collaboration, intuition, creativity, and experience. Verifying and interpreting tacit knowledge is challenging, particularly in terms of reliability and ethical considerations, as it is not derived from human personal experience, validated sensors, or deductive inferences, but from patterns in data. A semantics for RCKG tacit knowledge requires addressing the nature of reactive inferences in a continuous space with discrete input and output that can be context-sensitive.
This puts RCKG apart from most logics. For example, RCKG semantics is plausibility-preserving rather than truth-preserving. Its operators for entailment, necessity, possibility, etc. would be interpreted differently from traditional truth-theoretic ones, and the interaction between classical truth-preserving and plausibility-preserving axioms needs to be negotiated. E.g., plausibility does not guarantee truth, while truth guarantees plausibility, leading to invalidity of certain mixed transitivity chains. Also, tacitness is typically non-monotonic and context-dependent. Furthermore, multiple tacit operators emerge which can vary in their reasoning behaviour: social implicatures can have different logical properties (group relativity, role sensitivity, status dependency, intensity, valence, gradability, source dependency, etc.) from affective evocations, evaluative judgments, causal consequences, metaphorical blending, etc. However, being able to represent such a variety of tacit knowledge by using a unique continuous model, and forcing it to adhere to the valid logical axioms of an amodal KG is a big step forward to use logic in practical, large scale, and flexible contexts that depend on sophisticated implicit knowledge.
Constraining RCKGs to generate knowledge within precise boundaries poses significant challenges. However, we deem the effort with formalizing tacit knowledge valuable. This involves limiting the continuous geometric space to a discrete dimension, which allows, for instance, to get only the relevant knowledge with respect to a given objective (e.g., generating climate projections that harmonise experts’ opinions and target expected environmental scenarios). To address these issues we started formalizing and implementing a novel architecture called Logic Augmented Generation (LAG). LAG integrates RCKGs with Semantic KGs (SKGs) that serve as a discrete, heuristic layer for in-context learning within RCKG’s continuous knowledge space, enforcing hybrid logical consistency, factual boundaries, and fostering interoperability with existing SKGs in the Semantic Web. Figure 1 exemplifies the architecture of a LAG that is composed of: (i) an amodal SKG automatically gathered from text using FRED, representing user queries; (ii) an LLM that is prompted with the amodal SKG, additional prompt heuristics that can be used to refine the context incrementally, and the SKG scoping the desired semantics and providing existing factual knowledge; (iii) an extended SKG produced by LAG, which contains newly generated triples that comply with the SKG, and addresses the desired user queries.
Figure 1: The architecture of LAG.
A SKG is enriched as a RCKG via in-context learning that requires adequate prompt engineering strategies. The study of effective prompt engineering strategies to inject the logical and factual knowledge of a SKG into a RCKG is another challenge that LAG must address. Evolutions of Chain-of-Thought, such as Metacognitive Prompting [13], can be beneficial to LAG, as shown by preliminary results presented in [14].
An example that explains how LAG supports collective intelligence for medical diagnostics begins with the following simple medical case:
“A 38-year-old male presents with fever and a dry cough that has persisted for the last 4 days. He has no significant medical history, though he recently returned from a business trip.”
A possible extended KG generated by LAG is presented in Figure 2.
Figure 2: Example of possible extended KG generated by LAG.
This KG captures causal consequences emerging as tacit knowledge, i.e., the predicate caus:triggeringCauseFor, from the presented medical case. Additionally, Wikidata is used as an external SKG to perform knowledge reconciliation. In fact, the entities wd:Q61509, wd:Q38933, wd:Q35805 come from Wikidata and represent the concepts travel, fever, and cough, respectively. The extended KG makes formally emerge that a trip is the potential triggering cause for both fever and cough, a cause that was only tacitly assumed in the case description. Hence, experts might rely on the extended KG in order to provide their opinions that can be aggregated to the extended KG following the same process. This approach fosters collaboration, improves decision-making, and enhances interpretability and reliability in domains such as medical diagnostics and climate services, where precise, accountable, and actionable insights are critical.
5 Conclusions and future work
In this work we introduce Logic Augmented Generation (LAG), a novel solution that integrates Semantic Knowledge Graphs (SKGs) with Reactive Continuous Knowledge Graphs (RCKGs) to enforce logical consistency and blend symbolic and continuous reasoning. RCKGs are framed as a neuro-symbolic approach that uses Large Language Models (LLMs) to extract and represent knowledge, incorporating tacit insights and contextual nuances. This hybrid model enables context-aware knowledge generation and addresses challenges like formalising tacit knowledge, enhancing prompt engineering, and aligning plausibility- and truth-preserving logic. The potentiality of LAG is illustrated by using collective intelligence as a case study. Future work will study SKG-RCKG integration, advance prompt strategies, and expand operators for representing tacit semantics.
Acknowledgements
This work was supported by the European Union’s Horizon Europe research and innovation programme within the context of the project HACID (Hybrid Human Artificial Collective Intelligence in Open-Ended Domains, grant agreement No 101070588), and FAIR (Future Artificial Intelligence Research PNRR project, code PE00000013 CUP 53C22003630006)